OpenStudy (anonymous):

THE SUM OF FOUR CONSECUTIVE INTEGERS IS 286. FIND THE INTEGERS.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

x+(x+1)+(x+2)+(x+3)= 286

OpenStudy (anonymous):

X+(X+2)+(X+4)+(X+6)

OpenStudy (skullpatrol):

Do you know what "consecutive" integers means?

OpenStudy (anonymous):

x+1x+2x+3x=286??

OpenStudy (anonymous):

consecutive means back to back...in a row

OpenStudy (skullpatrol):

consecutive integers: Numbers obtained by counting by ones from any number in the set of integers.

OpenStudy (anonymous):

ok I will write that down.

OpenStudy (skullpatrol):

This is the set of integers: {...-4, -3, -2, -1, 0, 1, 2, 3, 4,...}

OpenStudy (anonymous):

I saved that info...

OpenStudy (anonymous):

x+1x+2x+3x=286??

OpenStudy (anonymous):